# Representation noising prevents harmful fine-tuning on LLMs

### v2

@@ Abstract
Releasing open-source large language models (LLMs) presents a dual-use risk since bad actors can easily fine-tune these models for harmful purposes. Even without the open release of weights, weight stealing and fine-tuning APIs make closed models vulnerable to harmful fine-tuning attacks (HFAs). While safety measures like preventing jailbreaks and improving safety guardrails are important, such measures can easily be reversed through fine-tuning. In this work, we propose Representation Noising (RepNoise), a defence mechanism that is effective even when attackers have access to the weights. RepNoise works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. Importantly, our defence is also able to generalize across different subsets of harm that have not been seen during the defence process as long as they are drawn from the same distribution of the attack set. Our method does not degrade the general capability of LLMs and retains the ability to train the model on harmless tasks. We provide empirical evidence that the effectiveness of our defence lies in its “depth”: the degree to which information about harmful representations is removed across all layers of the LLM.
@@ 1 Introduction
We propose Representation Noising (RepNoise) as the first effective defence against in-distribution harmful fine-tuning attacks (HFAs) for LLMs in the natural language generation setting where the defender has no control of the model after the attacker attains its weights. Our work is inspired by the observation that safety mechanisms in LLMs are concentrated in a small proportion of the model weights (identified through ablation studies in [70]) and displace rather than replace harmful capabilities (identified by probing studies in [39, 42]); despite showing safe behaviour at inference-time, harmful behaviour can be easily recovered [56]. RepNoise works by removing the information structure of harmful representations such that they are much harder to recover during subsequent HFAs (fig. 1). We refer readers to § A.1 for precise definitions of representations, information in representations, and removing information.
@@
    We present extensive experimental evidence that our method prevents training on harmful question-answering and toxic content generation tasks while maintaining the ability to train the model on harmless tasks and preserving LLM capability (§ 4).
@@ 2 Harmful Fine-tuning and Defence Criteria
Suppose an attacker has access to a set of model weights for a safety-aligned model—such as llama2-7b-chat [65]—which should refuse the attacker’s harmful requests. Let $M_{\theta[t=0]}$ indicate a model $M$ with the parameters $\theta$, where $t$ indicates the training steps taken during the HFA. The initial unattacked model is indexed at $t=0$. The attacker utilizes some harmful dataset $D_{\text{harmful}}$ to train a LLM to be harmful (i.e. minimize language modelling loss on this dataset). The defender considers this behaviour unacceptable and does not want their model to be able to be trained towards this end. $D_{\text{harmful}}$ consists of prompts $X=\{X_{i}\}^{n}_{i=1}$ and target responses $Y=\{Y_{i}\}^{n}_{i=1}$. Using the harmful dataset, a malicious actor attempts to find parameters at training step $t^{*}$ that minimizes eq. 1, resulting in a model that is able to behave in a way designated harmful by the defender:
@@
where $\mathcal{L}(M_{\theta}(X),Y)$ is a typical causal language modeling loss.
Harmful Fine-tuning Success: A HFA is formally designated a success if it causes the subject model to exceed a chosen threshold $\phi$ set by the defender on a given safety metric $g(\cdot):g(M_{\theta})>\phi$. Examples of safety metrics include Attack Success Rate [50, ASR] and probability outputs from the Jigsaw Perspective API [43]. In this formalism, a viable defence is one in which the attacker will have to spend more training steps $t$ to achieve $g(M_{\theta})>\phi$ than they can afford in time and/or compute.
@@
To defend against HFAs, Rosati et al. [58] provides four conditions for a successful defence, called Immunization Conditions (IC). We introduce them below in order to motivate our search for a method that fulfils them theoretically and experimentally.
@@
A model that fulfils these conditions is said to be “immunized”. § 4 provides an operationalization of these conditions. Below, we provide the a defence that fulfils all these conditions.
@@ 3 Method
We propose Representation Noising, a method which fulfils the Immunization Criteria by reducing the mutual information between intermediate activations of harmful text sequences (their representations) and harmful text inputs before the attacker gains access and performs an HFA. Recall that after the attacker has access, the defender cannot intervene. There is evidence that current safety methods only suppress or route around preserved harmful representations [70, 39, 42]. This allows harmful behaviour to be easily recovered through HFAs. RepNoise instead aims to remove information about harmful tasks from intermediate activations over harmful text squences, to make it difficult for the model to relearn such information in future. The formal definition of information removal removal is based on mutual information between intermediate activations and generative outputs based on those activations and is specified in § A.1 which we encourage review before proceeding. RepNoise consists of a three-part loss: (i) reduce the predictive information in the weights for generating harmful outputs, (ii) retain capabilities by minimizing loss on harmless inputs, and (iii) push harmful representations towards random noise to remove harmful information. Fig. 1 presents a high-level intuition of our method.
@@
Our goal is to derive a loss function which will minimize the likelihood of recovering the mutual information $I(Z_{harmful};Y_{harmful})$ which is a quantity that measure how effective intermediate activations (or representations) $Z_{harmful}$ of harmful input sequences $X_{harmful}$ are at predicting the output token distribution $Y_{harmful}$. We are motivated by the observation in [1] that the number of training steps taken to fine-tune a model trained on one source task to another target task is minimized by $M_{\theta[t^{*}]}$) can be modelled as a transition probability over paths (or training trajectories) in a loss landscape. Formally in the language modeling case, a task here is a token output distribution over some dataset $D$. The source task is our initial pre-training distribution and the target task is the generative distribution of harmful text tokens in $D_{\text{harmful}}$. The loss landscape is the space ($\mathbf{R^{|}\theta|}$) of the value of a loss function $\mathcal{L}_{D}$ at every value of $\theta$. Paths or training trajectories in this landscape are the sequence of parameter configurations $\theta_{i}$ during a training procedure.
@@
A full proof of this is given in appendix A. Based on information bottleneck theory [64], we view multiple-layer neural networks as consisting of an encoder which maps the inputs $X$ to representations $Z$ and a decoder which maps the learned $Z$ to outputs $Y$. From Wang et al. [69] theorem 2, we can minimize the mutual information $I(Z;Y)$ directly by performing gradient ascent, which would decrease both the static distance and the reachability condition eq. 2 (see appendix A). However, we only want to minimize the transition probability that minimizes loss on a harmful dataset $D_{\text{harmful}}$. Therefore we only want to consider minimizing the mutual information $I(Z_{harmful};Y_{{harmful}})$ where harmful text sequences $X_{harmful}$ are passed through the model to construct intermediate activations $Z_{harmful}$ which are subsequently used to generate the output tokens $Y_{harmful}$.
@@
Gradient ascent over harmful samples can degrade overall language modelling capabilities, so we add a term to ensure that performance over harmless samples is not degraded. In view of the immunization conditions in Section 2, this ensures stability. Combining them, we get an adversarial loss function:
@@
As we see later (Section 5), simply minimizing adversarial loss does not effectively remove the ability to predict harmful text sequences from the activations over harmful text sequences. This is because performing gradient ascent only minimizes $I(Y;Z)$—the mutual information between the output and decoder layers. However, according to the data processing inequality [71], $I(Y;X)\leq I(Z;X)$ despite having low mutual information $I(Y;Z)$, the mutual information between the inputs and the encoder layers $I(Z;X)$ can still be high. Consequently, it is possible for representations to retain the ability to generate harmful token sequences.
The above data processing inequality implies that minimizing $I(X;Z)$ minimizes $I(X;Y)$. We can do this by minimizing the KL divergence between the distribution of harmful representations given harmful input token sequences $\mathcal{P}_{\theta[t=0]}(Z\,|\,X)$ and Gaussian noise $\mathcal{N}(0,\textbf{{I}})$: if these two distributions were the same then $Z$ would have no information about $X$ [71]. This yields the noise loss $\ell_{\text{noise}}=KL(p(Z\,|\,X)\>||\>\mathcal{N}(0,\textbf{{I}}))$. Combining this with eq. 4 using another regularization term $\beta$, we get the loss function for Representation Noising (RepNoise) which satisfies Theorem 1.
@@ 4.1 Resistance
Table 1 shows results for the harmful QA task. HFAs without any defence mechanism substantially increase the harmfulness score (Base) on the base model. Attacks with higher learning rates and more data tend to be stronger. This replicates previous results about the effectiveness of HFAs at circumventing safety training in LLMs [56, 72, 8, 54, 44]. Evaluating our defence mechanisms, Security Vectors provides some resistance but RepNoise is the only defence method to consistently able to provide significant resistance across all attacks (Mann-Whitney $U$-test, $p<0.001$). [^2] We note that stronger attacks with thorough hyperparameter search can still succeed against RepNoise which we discuss in Appendix E.1.
@@ 4.3 Trainability
We further evaluated whether fine-tuning on a harmless task results in undoing safety guards or makes models more susceptible to HFAs. After fine-tuning on each GEM dataset, a HFA is performed with $3\times 10^{-5}$ with 1k samples from BeaverTails as above. Unlike the results of Qi et al. [56], both the base model and RepNoise are not made more harmful after harmless fine-tuning on GEM. However, training on GEM does seem to make the HFA more effective (readers can compare with the same attack in table 1). Even for RepNoise we see a small increase in attack efficacy after training the model on CACAPO which indicates the possibility that additional harmless fine-tuning could undo the RepNoise defence: a vulnerability which future work should explore.
We replicated Qi et al. [56] results in § E.3 (which RepNoise still provides an effective defence for) and the primary difference is that they are using a general instruction-following dataset rather than a specific task like structured generation to fine-tune the models.
@@ 4.4 Generalization
The BeaverTails dataset categorizes samples into 14 types of harm. We evaluate the generalization performance of RepNoise by withholding five categories of harm when performing the defence and then evaluate the attack by performing an attack using 1k samples from that subset. We also perform an additional experiment (Half) where RepNoise is applied using 5k randomly selected samples from BeaverTails and a subsequent attack is performed using 5k unseen samples.
@@
The results in Table 5 show that a defence using RepNoise is able to generalize to a defence against HFAs performed with unseen samples and unseen types of harm. However, it is important to note that these attacks are still in-distribution since the unseen types of harm are still drawn from the same BeaverTails distribution that the defender has seen. Importantly, RepNoise is not an effective defence against unseen sample out-of-distribution which we demonstrate in § E.2 using a distribution shift in the attack set to the HEX-PHI attack set [56].
@@ 5 Mechanistic Analysis
We conjecture that RepNoise is effective because it reduces the information about harmfulness in representations across all layers of the neural network, making them harder to recover. This is inspired by observations from previous studies, which found that popular safety methods merely route around the harmful representations [42], that fine-tuning only learns a wrapper on top of existing representations [39], and that harmful representations are easily recovered [70].
@@ 6 Related Work
Some prior work addresses the attenuation of safety fine-tuning’s influence on model behavior which typically occurs during benign fine-tuning. [78] achieve this for vLLMs with an instruction fine-tuning dataset which contains safety material and [49] do so for LLMs by modifying the prompt template used during fine-tuning. [32] use a modified LoRA algorithm for fine-tuning which maintains safety influence and [74] use a model fusion-based technique to get around the limitations of performing safety fine-tuning either before or after fine-tuning on tasks. Other solutions could benefit from methods that correct the general tendency for models to perform more poorly in some domains after being fine-tuned in others, such as the method presented in [51]. Though these methods may be sufficient for their stated goals, our work aims to prevent the effects of harmful fine-tuning, regardless of whether they come about from benign or harmful fine-tuning.
@@
Unlearning
While Unlearning is concerned with forgetting a particular sample, class, or concept, we focus on whole task domains. Some works do measure the ‘relearn time’ of an unlearned sample [63, 23], which is similar to Resistance. A few papers also relate the information content in the model weights to unlearning [23, 24], but are too expensive to apply to LLMs. Some sample unlearning methods also use noise to achieve unlearning [34, 63, 45], but cannot be applied to whole domains. We initially explored additional closed-form unlearning methods [17, 18], but they did not work for our natural language generation case since they resulted in complete model degradation.
Domain Authorization and Negative Transfer
Finally, our method can be seen as a specialized form of domain authorization [68, 69, 14] where a model provider wants to authorize a model for only a specific use either at inference or training time. While we focus on harmful uses, we believe that our method would be applicable to specialized training-time authorization on any domain. Negative Transfer studies poor training-time transfer across domains [76]. While research usually aims to reduce it, increasing negative transfer might inspire future immunization methods. Our theoretical approach relies on work in domain adaptation and transfer learning [1] which poses a transition probability that our loss function attempts to minimize. Continuations of this theoretical analysis could aim to develop methods with guarantees of resistance.
@@ 7 Limitations
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI) and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada and the Vector Institute. FR is supported by a Canada CIFAR Chair in AI. We thank Ben Rank for providing the implementation of Vaccine that we use in this paper. The project on which this report is based was funded by the Federal Ministry of Education and Research under the funding code 16KIS2012. The responsibility for the content of this publication lies with the author

### v4

# Representation Noising: A Defence Mechanism Against Harmful Finetuning

@@ Abstract
Releasing open-source large language models (LLMs) presents a dual-use risk since bad actors can easily fine-tune these models for harmful purposes. Even without the open release of weights, weight stealing and fine-tuning APIs make closed models vulnerable to harmful fine-tuning attacks (HFAs). While safety measures like preventing jailbreaks and improving safety guardrails are important, such measures can easily be reversed through fine-tuning. In this work, we propose Representation Noising (RepNoise), a defence mechanism that operates even when attackers have access to the weights. RepNoise works by removing information about harmful representations such that it is difficult to recover them during fine-tuning. Importantly, our defence is also able to generalize across different subsets of harm that have not been seen during the defence process as long as they are drawn from the same distribution of the attack set. Our method does not degrade the general capability of LLMs and retains the ability to train the model on harmless tasks. We provide empirical evidence that the efficacy of our defence lies in its “depth”: the degree to which information about harmful representations is removed across all layers of the LLM. We also find areas where RepNoise still remains ineffective and highlight how those limitations can inform future research.
@@ 1 Introduction
We propose Representation Noising (RepNoise) as the first defence to mitigate in-distribution harmful fine-tuning attacks (HFAs) for LLMs in the natural language generation setting where the defender has no control of the model after the attacker attains its weights. Our work is inspired by the observation that safety mechanisms in LLMs are concentrated in a small proportion of the model weights (identified through ablation studies in [64]) and displace rather than replace harmful capabilities (identified by probing studies in [35, 38]); despite showing safe behaviour at inference-time, harmful behaviour can be easily recovered [51]. RepNoise works by removing the information structure of harmful representations such that they are much harder to recover during subsequent HFAs (fig. 1). We refer readers to § A.1 for precise definitions of representations, information in representations, and removing information.
@@
    We present extensive experimental evidence that our method can mitigate training on harmful question-answering and toxic content generation tasks while maintaining the ability to train the model on harmless tasks and preserving LLM capability (§ 4).
@@ 2 Harmful Fine-tuning and Defence Criteria
Suppose an attacker has access to a set of model weights for a safety-aligned model—such as llama2-7b-chat [60]—which should refuse the attacker’s harmful requests. Let $M_{\theta[t=0]}$ indicate a model $M$ with the parameters $\theta$, where $t$ indicates the training steps taken during the HFA. The initial unattacked model is indexed at $t=0$. The attacker utilizes some harmful dataset $D_{\text{harmful}}$ to train a LLM to be harmful (i.e. minimize language modeling loss on this dataset). The defender considers this behaviour unacceptable and does not want their model to be able to be trained towards this end. $D_{\text{harmful}}$ consists of prompts $X=\{X_{i}\}^{n}_{i=1}$ and target responses $Y=\{Y_{i}\}^{n}_{i=1}$. Using the harmful dataset, a malicious actor attempts to find parameters at training step $t^{*}$ that minimizes eq. 1, resulting in a model that is able to behave in a way designated harmful by the defender:
@@
where the loss $\mathcal{L}(M_{\theta}(X),Y)$ is a typical causal language modeling loss.
Harmful Fine-tuning Success: A HFA is formally designated a success if it causes the subject model to exceed a chosen threshold $\phi$ set by the defender on a given safety metric $g(\cdot):g(M_{\theta})>\phi$. Examples of safety metrics include Attack Success Rate [45, ASR] and probability outputs from the Jigsaw Perspective API [39]. In this formalism, a viable defence is one in which the attacker will have to spend more training steps $t$ to achieve $g(M_{\theta})>\phi$ than they can afford in time and/or compute.
@@
To defend against HFAs, Rosati et al. [53] provides four conditions for a successful defence, called Immunization Conditions (IC). We introduce them below in order to motivate our search for a method that fulfills them theoretically and experimentally.
@@
A model that fulfills these conditions is said to be “immunized.” § 4 provides an operationalization of these conditions. Below, we provide the a defence that fulfills all these conditions.
@@ 3 Method
We propose Representation Noising, a method which fulfills the Immunization Criteria by reducing the mutual information between intermediate activations of harmful text sequences (their representations) and harmful text inputs before the attacker gains access and performs an HFA. Recall that after the attacker has access, the defender cannot intervene. There is evidence that current safety methods only suppress or route around preserved harmful representations [64, 35, 38]. This allows harmful behaviour to be easily recovered through HFAs. RepNoise instead aims to remove information about harmful tasks from intermediate activations over harmful text sequences, to make it difficult for the model to relearn such information in future. The formal definition of information removal removal is based on mutual information between intermediate activations and generative outputs based on those activations and is specified in § A.1 which we encourage review before proceeding. RepNoise consists of a three-part loss: (i) reduce the predictive information in the weights for generating harmful outputs, (ii) retain capabilities by minimizing loss on harmless inputs, and (iii) push harmful representations towards random noise to remove harmful information. Fig. 1 presents a high-level intuition of our method.
@@
Our goal is to derive a loss function which will minimize the likelihood of recovering the mutual information $I(Z_{harmful};Y_{harmful})$ which is a quantity that measures how effective intermediate activations (or representations) $Z_{harmful}$ of harmful input sequences $X_{harmful}$ are at predicting the output token distribution $Y_{harmful}$. We are motivated by the observation in [1] that the number of training steps taken to fine-tune a model trained on one source task to another target task minimized by $M_{\theta[t^{*}]}$ can be modeled as a transition probability over paths (or training trajectories) in a loss landscape. Formally in the language modeling case, a task here is a token output distribution over some dataset $D$. The source task is our initial pre-training distribution and the target task is the generative distribution of harmful text tokens in $D_{\text{harmful}}$. The loss landscape is the space ($\mathbf{R^{|\theta|}}$) of the value of a loss function $\mathcal{L}_{D}$ at every value of $\theta$. Paths or training trajectories in this landscape are the sequence of parameter configurations $\theta_{i}$ during a training procedure for the iterates $i$.
@@
A full proof of this is given in appendix A. Based on information bottleneck theory [59], we view multiple-layer neural networks as consisting of an encoder which maps the inputs $X$ to representations $Z$ and a decoder which maps the learned $Z$ to outputs $Y$. From Wang et al. [63] theorem 2, we can minimize the mutual information $I(Z;Y)$ directly by performing gradient ascent ($\ell_{\text{ascent}}$), which would decrease both the static distance and the reachability condition eq. 2 (see appendix A).
@@
However, gradient ascent over harmful samples can degrade overall language modeling capabilities, so we add a term to ensure that performance over harmless samples is not degraded. In view of the immunization conditions in Section 2, this ensures stability. Combining them, we get an adversarial loss function:
@@
As we see later (Section 5), simply minimizing adversarial loss does not effectively remove the ability to predict harmful text sequences from the activations over harmful text sequences. This is because despite having low mutual information $I(Y;Z)$, the mutual information between the inputs and the encoder layers $I(Z;X)$ can still be high. Consequently, it is possible for representations to retain the ability to generate harmful token sequences.
The data processing inequality [65], $I(Y;X)\leq I(Z;X)$ implies that minimizing $I(X;Z)$ minimizes $I(X;Y)$. We can do this by minimizing the KL divergence between the distribution of harmful representations given harmful input token sequences $\mathcal{P}_{\theta[t=0]}(Z\,|\,X)$ and Gaussian noise $\mathcal{N}(0,\textbf{{I}})$: if these two distributions were the same then $Z$ would have no information about $X$ [65]. This yields the noise loss $\ell_{\text{noise}}=KL(p(Z\,|\,X)\>||\>\mathcal{N}(0,\textbf{{I}}))$. Combining this with eq. 4 using another regularization term $\beta$, we get the loss function for Representation Noising (RepNoise) which satisfies Theorem 1.
@@ 4.1 Resistance
Table 1 shows results for the harmful QA task. HFAs without any defence mechanism substantially increase the harmfulness score (Base) on the base model. Attacks with higher learning rates and more data tend to be stronger. This replicates previous results about the effectiveness of HFAs at circumventing safety training in LLMs [51, 66, 8, 49, 40]. Evaluating our defence mechanisms, Security Vectors provides some resistance but RepNoise is the only defence method to consistently able to provide significant resistance across all attacks (Mann-Whitney $U$-test, $p<0.001$). [^2] We note that stronger attacks with thorough hyperparameter search can still succeed against RepNoise which we discuss in Appendix E.1 so it should not yet be considered as a comprehensive defence.
@@ 4.3 Trainability
We further evaluated whether fine-tuning on a harmless task results in undoing safety guards or makes models more susceptible to HFAs. After fine-tuning on each GEM dataset, a HFA is performed with $3\times 10^{-5}$ with 1k samples from BeaverTails as above. Unlike the results of Qi et al. [51], both the base model and RepNoise are not made more harmful after harmless fine-tuning on GEM. However, training on GEM does seem to make the HFA more effective (readers can compare with the same attack in table 1). Even for RepNoise we see a small increase in attack efficacy after training the model on CACAPO which indicates the possibility that additional harmless fine-tuning could undo the RepNoise defence, a vulnerability which future work should explore.
We replicated the benign and AOA attacks of Qi et al. [51] results in § E.3 and found that RepNoise can mitigate them.
@@ 4.4 Generalization
The BeaverTails dataset categorizes samples into 14 types of harm. We evaluate the generalization performance of RepNoise by withholding five categories of harm when performing the defence training and then evaluate the attack by performing an attack using 1k samples from that subset. We also perform an additional experiment (Half) where RepNoise is trained using 5k randomly selected samples from BeaverTails and a subsequent attack is performed using 5k unseen samples.
@@
The results in Table 5 show that a defence using RepNoise is able to generalize to a defence against HFAs performed with unseen samples and unseen types of harm. However, it is important to note that these attacks are still in-distribution since the unseen types of harm are still drawn from the same BeaverTails distribution that the defender has seen. Importantly, RepNoise is not an effective defence against unseen sample out-of-distribution which we demonstrate in § E.2 using a distribution shift in the attack set to the HEX-PHI attack set [51]. A comprehensive analysis of additional attacks (appendix E) and ablations (appendix J) are presented so readers can clearly understand the limitations of RepNoise.
@@ 5 Mechanistic Analysis
We conjecture that RepNoise works because it reduces the information about harmfulness in representations across all layers of the neural network, making them harder to recover. This is inspired by observations from previous studies, which found that popular safety methods merely route around the harmful representations [38], that fine-tuning only learns a wrapper on top of existing representations [35], and that harmful representations are easily recovered [64].
@@ 6 Related Work
Some prior work addresses the attenuation of safety fine-tuning’s influence on model behavior which typically occurs during benign fine-tuning. [71] achieve this for LLMs with an instruction fine-tuning dataset which contains safety material and [44] do so for LLMs by modifying the prompt template used during fine-tuning. [28] use a modified LoRA algorithm for fine-tuning which maintains safety influence and [68] use a model fusion-based technique to get around the limitations of performing safety fine-tuning either before or after fine-tuning on tasks. Other solutions could benefit from methods that correct the general tendency for models to perform more poorly in some domains after being fine-tuned in others, such as the method presented in [46]. Though these methods may be sufficient for their stated goals, our work aims to mitigate the effects of harmful fine-tuning, regardless of whether they come about from benign or harmful fine-tuning.
@@ 7 Limitations
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI) and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada and the Vector Institute. FR is supported by a Canada CIFAR Chair in AI. We thank Ben Rank for providing the implementation of Vaccine that we use in this paper. We’d also like to acknowledge the useful discussions and feedback we have had with other authors in the field namely Tiansheng Huang, Boyi Wei, and Xiangyu Qi. The project on which this report is based was funded by the Federal Ministry of Education and Research under the funding code 16KIS2012. The responsibility for the content of this publication lies with the author